# Primus: A Pioneering Collection of Open-Source Datasets for Cybersecurity LLM Training

### v2

@@ 2.2.1 Composition
We specifically collect cybersecurity-related text from diverse sources, including Blogs, News, Books, Websites, Wikipedia, and MITRE, guided by prior pretraining work Aghaei et al. (2022). For Blogs and News, we select content from government agencies, standards bodies, cybersecurity companies, media, and forums. Meanwhile, Books cover a wide range of cybersecurity topics, and we exclude covers, tables of contents, and appendices while treating each extracted page as a separate sample. We also collect Webpages from well-known cybersecurity companies, which may include product descriptions, company profiles, FAQs, and API documentation. In addition, Wikipedia does not provide a predefined cybersecurity subset, so we perform a custom filtering process. Each Wikipedia article is associated with one or more category tags, which can be further expanded into subcategory tags. Starting from the root category "Computer Security", we recursively traverse its subcategories, using GPT-4o to determine whether a category is cybersecurity-related [^3] The prompt is provided in the Appx.E (Fig.8). This process yields 375 relevant categories, from which we extract corresponding Wikipedia articles. For MITRE, we leverage obsidian-mitre-attack [^4] https://github.com/vincenzocaputo/obsidian-mitre-attack, which converts STIX data from the official repository into readable Markdown.
@@ 2.2.3 Augmentation
We find that some web-scraped data contains valuable information but suffers from poor readability due to irregular formatting, such as inconsistent line breaks. To address this, we adopt a rewriting approach inspired by Cosmopedia [^5] https://github.com/huggingface/cosmopedia, a reproduction of the high-quality synthetic dataset used in phi-1.5 Li et al. (2023b). Specifically, we prompt an LLM to rewrite the given text into a specific style, including blog posts, textbooks, and Q&A formats [^6] The prompt is provided in the Appx.E (Fig.9). To increase diversity, the rewriting LLM is randomly selected from GPT-4o, Llama-3.1-405B-Instruct, DBRX Mosaic (2024), and Claude 3.5 Sonnet Anthropic (2024).
@@ 2.3.1 Cybersecurity Classifier
We then use the classifier to score all FineWeb texts on a scale from 0 to 1, where higher scores indicate greater cybersecurity relevance. The distribution in Fig.3 shows that lower scores correspond to a significant increase in text volume. To determine an appropriate threshold for filtering, we first verify that whether texts with higher scores are truly cybersecurity-related. To do this, we leverage GPT-4o for accurate evaluation by dividing the scores into multiple bins, with dynamically adjusted bin sizes—smaller bins for lower scores—to account for the increased volume of data in lower score ranges. We randomly sample 50 texts from each bin and prompt GPT-4o [^7] The prompt is provided in the Appx.E (Fig.10) for classification. As shown in Fig.4, relevant text proportions remain above 60% at higher scores, but drop below 50% when scores fall below 0.003. Although incorporating some general text can help mitigate catastrophic forgetting Sun et al. (2019), we prioritize maintaining a majority of cybersecurity content. Therefore, we set the final threshold at 0.003, which corresponds to 15.3B of FineWeb data.
@@ 2.4 Primus-Instruct
After pre-training, we use Primus-Instruct for instruction fine-tuning to restore the instruction-following capability of the model. To achieve this, we design several hundred cybersecurity tasks covering common business scenarios, including explaining detected alerts, answering questions about retrieved security documents, analyzing executed suspicious commands, generating query languages for retrieving security events, and providing security recommendations and risk assessments for Terraform configurations. Each example is answered by GPT-4o, and we further use Claude 3.5 Sonnet as a judge [^8] The judge prompt is provided in the Appx.E (Fig.11) to discard samples with insufficiently helpful answers. In addition, we include several hundred multi-turn conversations on general topics generated by GPT-4o. As a result, these form Primus-Instruct, with statistics in Tab.3.
@@ 2.5 Primus-Reasoning
With the release of OpenAI’s reasoning model o1, an increasing number of studies have attempted to replicate its reasoning capabilities. One widely recognized approach is distillation, where reasoning samples with self-reflection from existing reasoning models are used to guide models in acquiring long-thought capabilities Huang et al. (2024); Liu et al. (2024). To this end, we select cybersecurity reasoning tasks from CTI-Bench [^9] A brief introduction to CTI-Bench is provided in Appx.C Alam et al. (2024) and prompt o1-preview one to two times per question to generate solutions with reasoning steps and reflection [^10] The prompt is provided in the Appx.E (Fig.12), applying rejection sampling to retain only the correctly answered samples. We also include DeepSeek-R1, obtained by directly querying its open-source model to access reasoning steps. The dataset statistics are shown in Tab.4.
@@ 3.2 Evaluation Settings
We follow the evaluation setup from the OpenAI technical report benchmarks with simple-eval [^11] https://github.com/openai/simple-evals, using a standardized prompt [^12] The prompt is provided in the Appx.E (Fig.13) that allows the model to articulate its reasoning before producing the final answer. Due to the formatting variability of CoT responses, we use GPT-4o-mini to extract the final answers before scoring.
@@ 4.1 Overview
In this section, we present the entire training pipeline, which consists of four key stages. First, we expand the model’s cybersecurity expertise and understanding through continued pre-training (Sec.4.2), which reinforces key cybersecurity concepts and enables the model to provide accurate information on security threats and mitigation strategies. Next, we restore its instruction-following capability through instruction fine-tuning (Sec.4.3), and further refine it through model merging to balance instruction-following and cybersecurity expertise. Finally, we train the model to develop reasoning capabilities on cybersecurity tasks (Sec.4.4) [^13] The training hyperparameters for each stage are provided in the Appx.D.
@@ 4.2 Pre-Training
+ Primus-Seed 0.7132 0.6608 0.6100 1.2848 0.2829 0.8600 0.4998 2.34$\uparrow$2.1%
+ Primus-FineWeb 0.7191 0.6600 0.6680 1.1499 0.3006 0.8620 0.4984 2.56$\uparrow$11.5%
+ Primus-Seed+FineWeb 0.7230 0.6676 0.6780 1.0912 0.3140 0.8660 0.5007 2.66$\uparrow$15.9%
@@
We train the model using the NeMo NVIDIA (2025) on four 8$\times$×H200 nodes, with training hyperparameters and details provided in Appx.D. To analyze the impact of different datasets, we conduct an ablation study by pre-training the model separately on each dataset and jointly on both for two epochs. The results in Tab.5 show that pre-training on either dataset improves the cybersecurity performance in the aggregate evaluation score. However, the largest improvement, 15.9%, is observed when pre-training on the combined dataset, so we adopt this model as the Llama-Primus-Base for subsequent training stages [^14] We also experimented with a 70B model in Q2 of Appx.A (FAQs).
@@ 4.3 Instruction Fine-Tuning and Merge
Llama-Primus-Instruct 0.7132 0.6660 0.6660 1.1161 0.3348 0.8640 0.4943 7.9063 4.21$\uparrow$2.4%
Llama-Primus-Merged 0.7191 0.6656 0.6620 1.1233 0.3387 0.8660 0.5062 8.2938 4.33$\uparrow$5.4%
Table 6: Performance comparison of Llama, the instruction-tuned Primus model, and their merge on cybersecurity and general benchmarks. The aggregated score (Agg.) is computed as 0.3×0.3$\times$0.3 × MT-Bench + 0.7×0.7$\times$0.7 × aggregated cybersecurity score (sum of all benchmarks except MT-Bench, with CTI-VSP negated due to the use of Mean Absolute Deviation, where lower is better). The best results are highlighted in bold.
While Llama-Primus-Base gains enhanced cybersecurity knowledge and understanding from pre-training, it tends to perform text completion rather than follow instructions. To address this, we further fine-tune it using the LLaMA-Factory Zheng et al. (2024) on 4$\times$×A100 GPUs for two epochs with Primus-Instruct (Sec.2.4), a carefully curated mixed dataset of cybersecurity tasks and general conversations, resulting in Llama-Primus-Instruct. In addition to the cybersecurity benchmarks, we also introduce MT-Bench Zheng et al. (2023), a multi-turn instruction-following evaluation benchmark spanning multiple domains using GPT-4 as a judge, which scores helpfulness on a scale of 1 to 10, allowing us to evaluate the overall instruction-following performance of the model. The results are shown in Tab.6, where the MT-Bench score and the aggregated cybersecurity benchmark score are further aggregated with a weight of 30/70 in the rightmost column.
Llama-Primus-Instruct maintains its advantage in cybersecurity while achieving an MT-Bench score of 7.91. However, this remains lower than the 8.35 of Llama, resulting in a limited improvement in the aggregated score (2.4%). To mitigate this, we apply DARE-TIES Yu et al. (2024); Yadav et al. (2023), a model merging technique that balances diverse capabilities—specifically, instruction-following and cybersecurity expertise in our case. We conduct a grid search over the merging ratio, setting Llama-Primus-Instruct:Llama-3.1-8B-Instruct to $(0.5+w)$:$(0.5-w)$ and varying $w$ from 0 to 0.5 in steps of 0.05. The optimal ratio that maximizes the aggregated score is found to be 0.75:0.25, with the merged model chosen as Llama-Primus-Merged. Notably, this configuration retains cybersecurity performance comparable to Llama-Primus-Instruct while restoring the MT-Bench to 8.29, almost equal to Llama, resulting in a 5.4% improvement in the aggregated score [^15] We provide more details in Q4 and Q5 of Appx.A (FAQs).
@@ 4.4 Reasoning Fine-Tuning
Llama-Primus-Merged 0.7191 $\uparrow$1.671
@@
Llama-3.1-8B-Instruct 0.7288 $\uparrow$3.03279.69
  + Distilled from o1-preview 0.7583 $\uparrow$7.21646.94
  + Distilled from DeepSeek-R1 0.7859 $\uparrow$11.11667.56
  + Distilled from (o1 + R1) 0.7780 $\uparrow$10.01615.54
Llama-Primus-Merged 0.7603 $\uparrow$7.49241.92
  + Distilled from o1-preview 0.7780 $\uparrow$10.0726.96
  + Distilled from DeepSeek-R1 0.8075 $\uparrow$14.21483.94
  + Distilled from (o1 + R1) 0.8193 $\uparrow$15.81467.40
@@ 6 Conclusion
We used Garak Derczynski et al. (2024), a toolkit that probes for hallucination, data leakage, prompt injection, misinformation, toxicity generation, jailbreaks, and many other vulnerabilities, to evaluate Llama-Primus-Merged. The results showed no significant differences compared to Llama (Appx.F). However, we still emphasize that the user is solely responsible for the content generated with the Primus model, as it lacks mechanisms to handle the disclosure of harmful, biased, or toxic content. Therefore, we strongly recommend that Primus be used for research purposes only. If used in production for natural language generation, users should independently assess the risks and implement appropriate safeguards.
@@ Appendix A FAQs
These details are provided in the appendix. The training hyperparameters are listed in Appx.D, and the prompts used for dataset construction are included in Appx.E.
@@
Yes, we extended our experiments to a 70B model by further pre-training Llama-3.1-Nemotron-70B-Instruct to obtain Llama-Primus-Nemotron-70B-Base. In addition to the dataset used for the 8B model, we supplemented its pre-training corpus with 7.6B tokens of cybersecurity content filtered from Nemotron-CC Su et al. (2024) (see Appx.B). The results in Tab.10 show an 11.2% gain in the aggregated cybersecurity benchmark score. We will also release this model under the MIT license. Due to its high computational cost, we did not conduct the dataset-combination ablation study on the 70B model that we performed on the 8B experiments.
@@
Table 11: Mean absolute error (MAE) between human expert scores and LLM scores across different Primus-Instruct tasks.
Yes, we conducted an experiment to measure the discrepancy between human experts and LLM judges under identical prompts. Specifically, in Sec.2.4 we used Claude 3.5 Sonnet to rate the helpfulness of responses in Primus-Instruct, discarding those that were not helpful enough [^17] The judge prompt is provided in the Appx.E (Fig.11). To validate Claude’s reliability as a judge, we randomly selected ten examples per task for human experts to score, then computed the differences between human, GPT-4o, and Claude ratings.
@@
Appendix B Primus-Nemotron-CC
We further extracted cybersecurity-related text from Nemotron-CC Su et al. (2024), which claims higher quality and more “unique” tokens than FineWeb (i.e., tokens remaining after global fuzzy deduplication). We scored each Nemotron-CC sample using the binary classifier trained in Sec.2.3 and partitioned the scores into multiple intervals. For each score interval, we sampled 1,000 examples, grouped them by length, sent them to GPT-4o-mini [^18] The prompt is provided in Appx.E (Fig.10) to verify whether they were truly cybersecurity-related, and then calculated the proportion of confirmed samples. The results are shown in Fig.6.
Figure 6: Ratio of cybersecurity-related text across different score bins in Nemotron-CC, grouped by sample length.
Figure 7: Ratio of cybersecurity-related text across score bins in the 1.0 ~ 0.9 range in Nemotron-CC.

### v3

@@ 2.2.1 Composition
We specifically collect cybersecurity-related text from diverse sources, including Blogs, News, Books, Websites, Wikipedia, and MITRE, guided by prior pretraining work Aghaei et al. (2022). For Blogs and News, we select content from government agencies, standards bodies, cybersecurity companies, media, and forums. Meanwhile, Books cover a wide range of cybersecurity topics, and we exclude covers, tables of contents, and appendices while treating each extracted page as a separate sample. We also collect Webpages from well-known cybersecurity companies, which may include product descriptions, company profiles, FAQs, and API documentation. In addition, Wikipedia does not provide a predefined cybersecurity subset, so we perform a custom filtering process. Each Wikipedia article is associated with one or more category tags, which can be further expanded into subcategory tags. Starting from the root category "Computer Security", we recursively traverse its subcategories, using GPT-4o to determine whether a category is cybersecurity-related [^3] The prompt is provided in the Appx.G (Fig.8). This process yields 375 relevant categories, from which we extract corresponding Wikipedia articles. For MITRE, we leverage obsidian-mitre-attack [^4] https://github.com/vincenzocaputo/obsidian-mitre-attack, which converts STIX data from the official repository into readable Markdown.
@@ 2.2.3 Augmentation
We find that some web-scraped data contains valuable information but suffers from poor readability due to irregular formatting, such as inconsistent line breaks. To address this, we adopt a rewriting approach inspired by Cosmopedia [^5] https://github.com/huggingface/cosmopedia, a reproduction of the high-quality synthetic dataset used in phi-1.5 Li et al. (2023b). Specifically, we prompt an LLM to rewrite the given text into a specific style, including blog posts, textbooks, and Q&A formats [^6] The prompt is provided in the Appx.G (Fig.9). To increase diversity, the rewriting LLM is randomly selected from GPT-4o, Llama-3.1-405B-Instruct, and DBRX Mosaic (2024).
@@ 2.3.1 Cybersecurity Classifier
We then use the classifier to score all FineWeb texts on a scale from 0 to 1, where higher scores indicate greater cybersecurity relevance. The distribution in Fig.3 shows that lower scores correspond to a significant increase in text volume. To determine an appropriate threshold for filtering, we first verify that whether texts with higher scores are truly cybersecurity-related. To do this, we leverage GPT-4o for accurate evaluation by dividing the scores into multiple bins, with dynamically adjusted bin sizes—smaller bins for lower scores—to account for the increased volume of data in lower score ranges. We randomly sample 50 texts from each bin and prompt GPT-4o [^7] The prompt is provided in the Appx.G (Fig.10) for classification. As shown in Fig.4, relevant text proportions remain above 60% at higher scores, but drop below 50% when scores fall below 0.003. Although incorporating some general text can help mitigate catastrophic forgetting Sun et al. (2019), we prioritize maintaining a majority of cybersecurity content. Therefore, we set the final threshold at 0.003, which corresponds to 15.3B of FineWeb data.
@@ 2.4 Primus-Instruct
After pre-training, we use Primus-Instruct for instruction fine-tuning to restore the instruction-following capability of the model. To achieve this, we design several hundred cybersecurity tasks covering common business scenarios, including explaining detected alerts, answering questions about retrieved security documents, analyzing executed suspicious commands, generating query languages for retrieving security events, and providing security recommendations and risk assessments for Terraform configurations. Each example is answered by GPT-4o, and we further use Claude 3.5 Sonnet Anthropic (2024) as a judge [^8] The judge prompt is provided in the Appx.G (Fig.11) to discard samples with insufficiently helpful answers. In addition, we include several hundred multi-turn conversations on general topics generated by GPT-4o. As a result, these form Primus-Instruct, with statistics in Tab.3.
@@ 2.5 Primus-Reasoning
With the release of OpenAI’s reasoning model o1, an increasing number of studies have attempted to replicate its reasoning capabilities. One widely recognized approach is distillation, where reasoning samples with self-reflection from existing reasoning models are used to guide models in acquiring long-thought capabilities Huang et al. (2024); Liu et al. (2024). To this end, we select cybersecurity reasoning tasks from CTI-Bench [^9] A brief introduction to CTI-Bench is provided in Appx.D Alam et al. (2024) and prompt o1-preview one to two times per question to generate solutions with reasoning steps and reflection [^10] The prompt is provided in the Appx.G (Fig.12), applying rejection sampling to retain only the correctly answered samples. We also include DeepSeek-R1, obtained by directly querying its open-source model to access reasoning steps. The dataset statistics are shown in Tab.4.
@@ 3.2 Evaluation Settings
We follow the evaluation setup from the OpenAI technical report benchmarks with simple-eval [^11] https://github.com/openai/simple-evals, using a standardized prompt [^12] The prompt is provided in the Appx.G (Fig.13) that allows the model to articulate its reasoning before producing the final answer. Due to the formatting variability of CoT responses, we use GPT-4o-mini to extract the final answers before scoring.
@@ 4.1 Overview
In this section, we present the entire training pipeline, which consists of four key stages. First, we expand the model’s cybersecurity expertise and understanding through continued pre-training (Sec.4.2), which reinforces key cybersecurity concepts and enables the model to provide accurate information on security threats and mitigation strategies. Next, we restore its instruction-following capability through instruction fine-tuning (Sec.4.3), and further refine it through model merging to balance instruction-following and cybersecurity expertise. Finally, we train the model to develop reasoning capabilities on cybersecurity tasks (Sec.4.4) [^13] The training hyperparameters for each stage are provided in the Appx.E.
@@ 4.2 Pre-Training
[…]
@@
[…]
@@ 4.3 Instruction Fine-Tuning and Merge
[…]
Table 6: Performance comparison of Llama, the instruction-tuned Primus model, and their merge on cybersecurity and general benchmarks. The aggregated score (Agg.) is computed as $0.3\times$ MT-Bench + $0.7\times$ aggregated cybersecurity score (sum of all benchmarks except MT-Bench, with CTI-VSP negated due to the use of Mean Absolute Deviation, where lower is better). The best results are highlighted in bold.
[…]
Llama-Primus-Instruct maintains its advantage in cybersecurity while achieving an MT-Bench score of 7.91. However, this remains lower than the 8.35 of Llama, resulting in a limited improvement in the aggregated score (2.4%). To mitigate this, we apply DARE-TIES Yu et al. (2024); Yadav et al. (2023), a model merging technique that balances diverse capabilities—specifically, instruction-following and cybersecurity expertise in our case. We conduct a grid search over the merging ratio, setting Llama-Primus-Instruct:Llama-3.1-8B-Instruct to (0.5+w)(0.5+w):(0.5w)(0.5-w) and varying $w$ from 0 to 0.5 in steps of 0.05. The optimal ratio that maximizes the aggregated score is found to be 0.75:0.25, with the merged model chosen as Llama-Primus-Merged. Notably, this configuration retains cybersecurity performance comparable to Llama-Primus-Instruct while restoring the MT-Bench to 8.29, almost equal to Llama, resulting in a 5.4% improvement in the aggregated score [^15] We provide more details in Q4 and Q5 of Appx.A (FAQs).
@@ 4.4 Reasoning Fine-Tuning
Llama-Primus-Merged 0.7191 $\uparrow$1.67% 1
@@
Llama-3.1-8B-Instruct 0.7288 $\uparrow$3.03% 279.69
  + Distilled from o1-preview 0.7583 $\uparrow$7.21% 646.94
  + Distilled from DeepSeek-R1 0.7859 $\uparrow$11.1% 1667.56
  + Distilled from (o1 + R1) 0.7780 $\uparrow$10.0% 1615.54
Llama-Primus-Merged 0.7603 $\uparrow$7.49% 241.92
  + Distilled from o1-preview 0.7780 $\uparrow$10.0% 726.96
  + Distilled from DeepSeek-R1 0.8075 $\uparrow$14.2% 1483.94
  + Distilled from (o1 + R1) 0.8193 $\uparrow$15.8% 1467.40
@@ 6 Conclusion
We used Garak Derczynski et al. (2024), a toolkit that probes for hallucination, data leakage, prompt injection, misinformation, toxicity generation, jailbreaks, and many other vulnerabilities, to evaluate Llama-Primus-Merged. The results showed no significant differences compared to Llama (Appx.H). However, we still emphasize that the user is solely responsible for the content generated with the Primus model, as it lacks mechanisms to handle the disclosure of harmful, biased, or toxic content. Therefore, we strongly recommend that Primus be used for research purposes only. If used in production for natural language generation, users should independently assess the risks and implement appropriate safeguards.
@@ Appendix A FAQs
These details are provided in the appendix. The training hyperparameters are listed in Appx.E, and the prompts used for dataset construction are included in Appx.G.
@@
Yes, we extended our experiments to a 70B model by further pre-training Llama-3.1-Nemotron-70B-Instruct to obtain Llama-Primus-Nemotron-70B-Base. In addition to the dataset used for the 8B model, we supplemented its pre-training corpus with 7.6B tokens of cybersecurity content filtered from Nemotron-CC Su et al. (2024) (see Appx.C). The results in Tab.10 show an 11.2% gain in the aggregated cybersecurity benchmark score. We will also release this model under the MIT license. Due to its high computational cost, we did not conduct the dataset-combination ablation study on the 70B model that we performed on the 8B experiments.
@@
Table 11: Mean absolute error (MAE) between human expert scores and LLM scores across different Primus‐Instruct tasks.
Yes, we conducted an experiment to measure the discrepancy between human experts and LLM judges under identical prompts. Specifically, in Sec.2.4 we used Claude 3.5 Sonnet to rate the helpfulness of responses in Primus-Instruct, discarding those that were not helpful enough [^17] The judge prompt is provided in the Appx.G (Fig.11). To validate Claude’s reliability as a judge, we randomly selected ten examples per task for human experts to score, then computed the differences between human, GPT-4o, and Claude ratings.
@@
• Q7: What are the structures of the datasets proposed in this paper?
The schema for each dataset is provided in Appx.B, including the fields it contains, their descriptions, and license information.
• Q8: Could you provide some sample responses from Llama-Primus that demonstrate its capabilities?
In Appx.F, we present and compare the responses of Llama-Primus-Merged and Llama-3.1-8B-Instruct on a question selected from the CTI-MCQ dataset.
• Q9: Could the improvement in CISSP scores after training on Primus-Reasoning be attributed to the inclusion of CISSP answers, thereby leading to potential data leakage?
To ensure the rigor of our experiments, we applied the standard N-gram decontamination method from EleutherAI’s llm-eval-harness to identify any overlapping content, following the approach described in the GPT-3 paper Brown et al. (2020) (with N set to 13 by default). Specifically, we concatenated each CISSP question and answer pair, and likewise concatenated the message contents of each sample in Primus-Reasoning, then generated N-grams for both and checked for duplicates. The results are shown in Tab.15.
We found only 5 overlapping samples between Primus-Reasoning and the CISSP benchmark when lowering N to 8. However, manual inspection revealed that these overlaps were limited to generic question stems such as "Which of the following best describes how an" and "Which of the following is an example of," rather than actual cybersecurity concepts or substantive content. Therefore, we believe that potential data leakage is negligible.
N-gram Overlap Count
13 0
12 0
11 0
10 0
9 0
8 5
Table 15: Counts of overlapping N-grams between Primus-Reasoning and CISSP.
Appendix B Dataset Details
Fields.
All datasets of Primus-Pretraining (Primus-Seed, Primus-FineWeb, and Primus-Nemotron-$C$) have the same structure and set of fields, as shown in Tab.16. Similarly, Primus-Instruct and Primus-Reasoning have a unified schema, which is detailed in Tab.17.
License.
All datasets proposed in this paper are released under the ODC-BY license. Additionally, compliance with the Terms of Use (ToU) or licenses of the original content sources is required. Some datasets are derived from existing ones. For example, Primus-FineWeb originates from FineWeb, and Primus-Nemotron-CC stems from Nemotron-CC. Both of these datasets are in turn based on Common Crawl, which requires compliance with its ToU. The Common Crawl ToU also requires adherence to the ToU of the original content owners.
As indicated in the field descriptions, all datasets of Primus-Pretraining include a url field that points to the original content source. We expect users to also respect the ToU or licenses of the original content providers.
Field Description
url The original source URL link corresponding to the sample.
source A coarse category of the sample’s source, such as Wikipedia or MITRE.
content The textual content of the sample.
time The crawling time of the sample, recorded in ISO 8601 format (e.g., 2024-12-31T00:00:00). For Primus-FineWeb and Primus-Nemotron-CC, only the year is recorded; to maintain format consistency, we append -12-31T00:00:00 after the year.
Table 16: Fields contained in each sample of Primus-Seed, Primus-FineWeb, and Primus-Nemotron-$C$.
Field Description
messages The conversation history stored in an alternating user/assistant format, e.g., [{"role": "user", "content": "..." }, {"role": "assistant", "content": "..." }, ... ].
prompt The first prompt from the user, i.e., the content of the first messages entry.
prompt_id A unique identifier for the sample.
Table 17: Fields contained in each sample of Primus-Instruct and Primus-Reasoning.
Appendix C Primus-Nemotron-CC
We further extracted cybersecurity-related text from Nemotron-CC Su et al. (2024), which claims higher quality and more “unique” tokens than FineWeb (i.e., tokens remaining after global fuzzy deduplication). We scored each Nemotron-CC sample using the binary classifier trained in Sec.2.3 and partitioned the scores into multiple intervals. For each score interval, we sampled 1,000 examples, grouped them by length, sent them to GPT-4o-mini [^18] The prompt is provided in Appx.G (Fig.10) to verify whether they were truly cybersecurity-related, and then calculated the proportion of confirmed samples. The results are shown in Fig.6.
Figure 6: Ratio of cybersecurity-related text across different score bins in Nemotron-$C$, grouped by sample length.
Figure 7: Ratio of cybersecurity-related text across score bins in the 1.0 ~ 0.9 range in Nemotron-$C$.